HUBBLE SPACE TELESCOPE DAILY REPORT #903 PERIOD COVERED: 0800 EST Daily Status Report as of 159/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed WFC 5010 (Determination of the Extragalactic distance Scale. III. M-101) The Wide Field Camera was used to make two images of the large Sc class spiral galaxy NGC-5457 in Ursa Major (aka, M-101) to look for Cepheid variable stars. Both images were received and displayed in the OSS. They both were reported to look okay. 1.2 Completed FOC/96 4804 (Ultraviolet Imaging of Nearby Galaxies FOC/96) The Faint Object Camera (f/96) was used to make one ultraviolet snapshot image of nearby galaxy UGC-3522. The exposure lasted ten minutes, and used filter F220W. The observations was received and displayed. It was reported to look okay. 1.3 Completed PC 3982 (Physical conditions in the Narrow - Lined Region) The Planetary Camera was used to make six observations of the active galactic region of Seyfert galaxy NGC-3393 to learn more about the structure of its nucleus. The first four observations were received and displayed in real time. They were reported to be okay. The last two observations occurred during LOS, and are still of the science tape recorder. They will be analyzed later. 1.4 Completed GHRS/2 4927 (GHRS SSA Peakup test - Part 1) The High Resolution Spectrograph was used to conduct a Small Science Aperture Peakup Test. The test involved using the return-to-brightnest algorithm with a small spiral search. The test consisted of a Large Science aperture onboard acquisition on Chi Lupi (aka,BD 75D325), a spectrum take in the Large Science Aperture to define the stellar flux, a slew to the Small Science Aperture followed by a new Peakup, and a mapping of the Small Science Aperture. The observations were executed as scheduled, and no problems were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 4 Successful : 4 Scheduled Reacquisitions : 10 Successful : 10 Losses of Lock : 0 2.2 FHST Updates: Scheduled : 10 Successful : 10 2.3 Significant HSTARs: None 2.4 Operations Note: A simulation of EVAs 1 and 2 were successfully conducted yesterday (Tuesday, June 7) with 24 TSARs being generated. Major simulated anomalies included the RSU2 gyros failing to power on and a SDM SA retraction failure. Each of these cases were referred to a management tiger team which recommended recovery scenarios. A debriefing will be conducted today at 3:00 P.M. 3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD: 3.1 Complete FOC/96 4353 (Hubble space Telescope: Imaging of Io: Cycle 3 High) 3.2 Complete Three Sets of WFC 4713 (WF/PC Cycle Two Calibration: Internal Delta Flats) 3.3 Complete FOS/RD 4772 (Black Holes, Stellar Dynamics and Populations in Nuclei) 3.4 Complete WFC 4713 (WF/PC Cycle Two Calibration: WFC Internal Delta Flats) 3.5 Complete WFC 4381 (HST Medium-Deep Survey - Cycle 3) 3.6 Complete WFC/PC 4109 (WF/PC Cycle 2 Calibration: Rapid Internal Monitor) 3.7 Complete PC 4140 (WF/PC Cycle 2 Calibration: Rapid Internal Monitor - Part 2) 3.8 Complete WFC 4929 (WF/PC Cycle 2 : Improved Dark Calibration) 3.9 Uplink Updated Ephemeris Tables, Per MOSES 2871 3.10 Complete FOC/96 3804 (Imaging of UV Bright Stars In Metal Rich Globular Clusters) 3.11 Complete PC 4902 (HST Medium-Deep Survey - Part 2: Low-Lat - Cycle 3 High) 3.12 Complete WFC 4713 (WF/PC Cycle 2 Calibration: WFC Internal Delta Flats) 3.13 Complete FOC/96 4804 (UV Imaging of Nearby Galaxies FOC/96) 4.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: None 5.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH -END-